# Universal Time-Series Representation Learning: A Survey

### v1

@@ 5. Data-Centric Approaches
In this group, we categorize the methods that focus on finding a new way to enhance the usefulness of the training data at hand. To capture the underlying patterns, trends, and relevant features within the time series, data-centric approaches prioritize engineering the data itself rather than focusing on model architecture and loss function design. We categorize data-centric approaches into three techniques: data augmentation, decomposition and transformation, and sample selection.
5.1. Data Augmentation Strategies
The studies in this category aim to increase the size and diversity of the training data. Especially in representation learning, most contrastive learning methods generate positive and negative samples by data augmentation. Also, different from other data types, time-series data augmentation should consider the unique properties in the time series, such as temporal dependencies, multi-scale dependencies, and inter-variable dependencies. We further categorize each method into random or policy-based augmentation.
5.1.1. Random Augmentation
In addition to hierarchical contrasting loss, TS2Vec (Yue et al., 2022) randomly treats two overlapping time segments as positive pairs for contrastive learning. Specifically, timestamp masking and random cropping are applied to the input time series to generate a context. Here, the contrasting objective in TS2Vec is based on the augmented context views, meaning that representations of the same sub-series in two augmented contexts should be consistent. As a result, we can obtain a robust contextual representation for each sub-series without introducing unappreciated inductive biases like transformation- and cropping-invariance, which proven effective for a wide range of downstream tasks. By considering frequency information, TF-C (Zhang et al., 2022b) aims to use both time and frequency features of time series to improve the representation quality. TF-C is claimed to be the first study to develop frequency-based contrastive augmentation exploiting spectral information and to explore time-frequency consistency in time series. It consists of a set of novel augmentations based on the characteristic of the frequency spectrum and produces a frequency-based embedding through contrastive instance discrimination. More specifically, TF-C introduces frequency-domain augmentation by randomly adding or removing frequency components, thereby exposing the model to a range of frequency variations. Recently, TS-CoT (Zhang et al., 2023e) facilitates the creation of diverse views for contrastive learning by enhancing robustness to noisy time series that contributes to the overall effectiveness of the representation learning.
5.1.2. Policy-based Augmentation
Instead of relying on randomness, several studies devised a specific criterion for data augmentation. TimeCLR (Yang et al., 2022a) addresses the challenges of applying self-supervised learning to time series by introducing dynamic time warping-based augmentation, which is designed to induce phase shifts and amplitude changes in time-series data while preserving its structural and feature information. BTSF (Yang and Hong, 2022) uses the entire time series as input and applies a standard dropout as an instance-level augmentation to generate different views for representation learning of time series. This paper argues that using dropout as the augmentation method can preserve the global temporal information and capture long-term dependencies of time series more effectively. The construction of contrastive pairs also ensures that the augmented time series do not change their raw properties, effectively reducing the potential false negatives and positives.
Luo et al. (2023) propose InfoTS, a meta-learning-based data augmentation method. InfoTS is an adaptive data augmentation method that uses information-aware criteria for selecting the optimal data augmentations to generate feasible positive samples for contrastive learning. The positive samples generated by InfoTS exhibit both high fidelity and diversity, which are essential attributes for contrastive learning. RIM (Aboussalah et al., 2023) is another new time-series augmentation method that generates more samples by using a recursive interpolation function from the original time series to be used in training. This method can control how much the augmented time series trajectory deviates from the original trajectory. Based on consistency regularization, Shin et al. (2023) propose a novel method of data augmentation called context-attached augmentation, which adds preceding and succeeding instances to a target instance to form its augmented instance, to fully exploit the sequential characteristic of time series in consistency regularization. This context-attached augmentation generates instances augmented with varying contexts while maintaining the target instance, thus avoiding direct perturbations on the target instance’s attributes. Another recent method that uses frequency-domain information is proposed by Demirel and Holz (2023). This method is based on a mixup technique for non-stationary quasi-periodic time series. This method aims to connect intra-class samples together to find order in the latent space. With a mixture of the magnitude and phase of frequency components together with tailored operations, it can prevent destructive interference of mixup.
For video data, Chen et al. (2022) use overlap augmentation techniques by sampling two subsequences with an overlap for each video. Overlapped timestamps are positives, and clips from other videos are negatives. Two timestamps neighboring each other also become positive pairs with Gaussian weight proportional to temporal distance. Zhang and Crandall (2022) introduce a novel self-supervised video representation learning technique emphasizing spatial and temporal features through a contrastive approach. Augmentation is used as a regularization tool, guiding the network to learn desired semantics in contrastive learning. FreqAug (Kim et al., 2023c) pushes the model to focus more on dynamic features via dropping spatial or temporal low-frequency components. For various downstream tasks, FreqAug stochastically eliminates particular frequency components from the video to help the latent representation better capture meaningful features from the remaining data. Kim et al. (2023b) propose DynaAugment by dynamically changing the augmentation magnitude over time in order to learn the temporal variations of a video. By Fourier sampling the magnitudes, it regulates the augmentations diversely while maintaining the temporal consistency.
Remark. Recent advancements in time-series representation learning have emphasized the importance of augmentation strategies tailored to capture both temporal and frequency characteristics effectively. While random augmentation methods leverage augmented context views for robust representation, policy-based approaches employ specific criteria to generate diverse and high-fidelity positive samples for contrastive learning. These augmentation techniques not only enhance the model’s ability to handle noisy and diverse time-series data but also ensure the preservation of essential temporal and spectral attributes.
5.2. Decomposition and Transformation
Unlike the above technique that increases the number of samples, methods in this subcategory aim to extract more meaningful information from the training data.
@@ 5.2.1. Time-Series Decomposition
To enhance video-level tasks, including video classification, and more detailed tasks like action segmentation, Behrmann et al. (2021) separate the representation space into stationary and non-stationary characteristics through contrastive learning from long and short views. Zeng et al. (2021) improve generalization across various downstream tasks by learning spatially-local/temporally-global and spatially-global/temporally-local features from audio-visual modalities to capture global and local information inside a video. This method enables the capturing of both slowly changing patch-level information and fast changing frame-level information. More recently, Yang et al. (2022b) propose a unified framework for joint audio-visual speech recognition and synthesis. Each modality is decomposed into modality-specific and modality-invariant features. Modality-invariant codebook enhances the alignment between the linguistic representation space of visual and audio modalities.
@@ 5.2.2. Input Space Transformation
To make effectively exploit vision models, Anand and Nayak (2021) transform 1D time series into 2D images and use the transformed images as features for the subsequent learning phase with models pretrained on large image datasets, such as ResNet50. Similarly, TimesNet (Wu et al., 2023) learns the temporal variations in the 2D space to enhance the representation capability by capturing variations both within individual periods and across multiple periods in the time series. It analyzes the time series from a new dimension of multi-periodicity by extending the analysis of temporal variations from 1D into 2D space. Besides, by transforming the 1D time series into a set of 2D tensors, TimesNet breaks the bottleneck of representation capability in the original 1D space, enabling well-performing vision backbones applicable to the transformed time series.
@@
To address tasks involving irregular time series data, some input space transformation techniques have been proposed. SplineNet (Biloš et al., 2022) generates splines from the input time series and directly utilizes them as input for a neural network. This approach introduces a learnable spline kernel to effectively process the input spline. MIAM (Lee et al., 2022b) considers multiple views of the input data, including time intervals, missing data indicators, and observation values. These transformed input data are then processed by the multi-view integration attention module to solve downstream tasks. Recently, Zhong et al. (2023) propose a novel multi-scale temporal patching approach to divide the input time series into non-overlapping patches along the temporal dimension in each layer. It treats the time series as patches. Different layers have different patch sizes such that they can focus on different time scales, making the representations more expressive.
For temporal graph networks, Chanpuriya et al. (2023) construct time-decayed line graph to train edge representation by creating a sparse matrix of proximities between temporal edges.
Remark. Time-series decomposition and input space transformation have significantly enhanced the capabilities of deep models for analyzing and processing time-series and video data. Notably, innovations in transforming 1D time series into 2D images or tensors, as well as methods addressing irregular time series, have expanded the applicability and performance of vision backbones.
5.3. Methods for Sample Selection
Methods in this group aim to generate or select the most useful samples based on the available training data for enhancing the informativeness of the learned representations.
5.3.1. Generative Methods

### v3

@@ 3. Data-Centric Approaches
This section presents the methods that focus on finding a new way to enhance the usefulness of the training data at hand. These approaches prioritize engineering the data itself rather than focusing on model architecture and loss function design to capture the underlying patterns, trends, and relevant features within the time series. As in Fig. 5, we categorize these data-centric approaches into two groups based on their objectives: improving data quality or increasing data quantity.
3.1. Improving Data Quality
3.1.1. Sample Selection
Sample selection aims to effectively choose the best samples from available training data for a particular learning scenario. An early study adopting sample selection for time-series representation learning is proposed by Franceschi et al. (2019). This work uses time-based negative sampling which determines several negative samples by independently choosing sub-series from other time series at random, whereas a sub-series within the referenced time series is consider a positive sample. This technique encourages representations of the input time segment and its sampled sub-series to be close to each other. MTRL (Chen et al., 2021a) utilizes discriminative samples to design a distance-weighted sampling strategy for achieving high convergence speed and accuracy.
@@ 3.1.2. Time-Series Decomposition
To enhance video-level tasks, including video classification, and more detailed tasks like action segmentation, Behrmann et al. (2021) separate the representation space into stationary and non-stationary characteristics through contrastive learning from long and short views. Zeng et al. (2021) improve generalization across various downstream tasks by learning spatially-local/temporally-global and spatially-global/temporally-local features from audio-visual modalities to capture global and local information inside a video. This method enables the capturing of both slowly changing patch-level information and fast changing frame-level information. More recently, Yang et al. (2022a) propose a unified framework for joint audio-visual speech recognition and synthesis. Each modality is decomposed into modality-specific and modality-invariant features. Modality-invariant codebook enhances the alignment between the linguistic representation space of visual and audio modalities.
@@ 3.1.3. Input Space Transformation
To effectively exploit vision models, DeLTa (Anand and Nayak, 2021) transforms 1D time series into 2D images and use the transformed images as features for the subsequent learning phase with models pretrained on large image datasets, such as ResNet50. Similarly, TimesNet (Wu et al., 2023) learns the temporal variations in the 2D space to enhance the representation capability by capturing variations both within individual periods and across multiple periods in the time series. It analyzes the time series from a new dimension of multi-periodicity by extending the analysis of temporal variations from 1D into 2D space. By transforming the 1D time series into a set of 2D tensors, TimesNet breaks the bottleneck of representation capability in the original 1D space, enabling well-performing vision backbones applicable to the transformed time series. Recently, Zhong et al. (2023) propose a novel multi-scale temporal patching approach to divide the input time series into non-overlapping patches along the temporal dimension in each layer. It treats the time series as patches. FITS (Xu et al., 2024) conducts interpolation in the frequency domain to extend time series and incorporates a low-pass filter to ensure a compact representation while preserving essential information.
@@
To address tasks involving irregular time series, some input space transformation techniques have been proposed. SplineNet (Biloš et al., 2022) generates splines from the input time series and directly utilizes them as input for a neural network. This approach introduces a learnable spline kernel to effectively process the input spline. MIAM (Lee et al., 2022a) considers multiple views of the input data, including time intervals, missing data indicators, and observation values. These transformed input data are then processed by the multi-view integration attention module to solve downstream tasks.
3.2. Increasing Data Quantity
3.2.1. Data Augmentation
Unlike other data types, augmenting time-series data requires special attention to its unique characteristics, such as temporal dependencies, multi-scale patterns, and inter-variable relationships. We categorize each method as either random or policy-based augmentation.
Random Augmentation
TS2Vec (Yue et al., 2022) randomly treats two overlapping time segments as positive pairs for contrastive learning. Specifically, timestamp masking and random cropping are applied to the input time series to generate a context. Here, the contrasting objective in TS2Vec is based on the augmented context views, meaning that representations of the same sub-series in two augmented contexts should be consistent. As a result, we can obtain a robust contextual representation for each sub-series without introducing unappreciated inductive biases like transformation- and cropping-invariance, which proven effective for a wide range of downstream tasks. By considering frequency information, TF-C (Zhang et al., 2022b) aims to use both time and frequency features of time series to improve the representation quality. It exploits spectral information and explores time-frequency consistency in time series with a set of novel augmentations based on the characteristic of the frequency spectrum. Specifically, TF-C introduces frequency-domain augmentation by randomly adding or removing frequency components, thereby exposing the model to a range of frequency variations. TimesURL (Liu and Chen, 2023) uses a frequency-temporal-based augmentation that generates augmented contexts by frequency mixing and random cropping to maintain the temporal property unchanged. Recently, TS-CoT (Zhang et al., 2023f) creates diverse views for contrastive learning by enhancing robustness to noisy time series that contributes to the overall effectiveness of the representation learning.
Policy-based Augmentation
Instead of relying on randomness, several studies devise a specific criterion for data augmentation. TimeCLR (Yang et al., 2022b) introduces a new data augmentation method based on dynamic time warping (DTW) that induces phase shifts and amplitude changes in time-series data while preserving its structural and feature information. Similarly, Yang et al. (2023a) introduce TempCLR, a contrastive learning framework for exploring temporal dynamics in video-paragraph alignment, presenting a new negative sampling strategy based on temporal granularity. TempCLR focuses on sequence-level comparison, using DTW to model the temporal order and dynamics of the data. BTSF (Yang and Hong, 2022) uses the entire time series as input and applies a standard dropout as an instance-level augmentation to generate different views for representation learning of time series. This paper argues that using dropout as the augmentation method can preserve the global temporal information and capture long-term dependencies of time series more effectively. The construction of contrastive pairs also ensures that the augmented time series do not change their raw properties, effectively reducing the potential false negatives and positives. Based on consistency regularization, Shin et al. (2023) propose a novel method of data augmentation called context-attached augmentation, which adds preceding and succeeding instances to a target instance to form its augmented instance, to fully exploit the sequential characteristic of time series in consistency regularization. This context-attached augmentation generates instances augmented with varying contexts while maintaining the target instance, thus avoiding direct perturbations on the target instance’s attributes. RIM (Aboussalah et al., 2023) is another new time-series augmentation method that generates more samples by using a recursive interpolation function from the original time series to be used in training. This method can control how much the augmented time series trajectory deviates from the original trajectory.
Using frequency-domain information, Demirel and Holz (2023) propose a novel data augmentation technique. This method is based on a mixup technique for non-stationary quasi-periodic time series that aims to connect intra-class samples together to find order in the latent space. With a mixture of the magnitude and phase of frequency components together with tailored operations, it can prevent destructive interference of mixup. MF-CLR (Duan et al., 2024) is a multi-frequency based method designed for self-supervised learning on multi-dimensional time series with varying frequencies. It uses a new data augmentation method called Dual Twister to enhance learning across different frequency levels by adding noise along both dimensions of an input time series to generate positive views with similar semantic meanings while preserving the original frequency structure.
Given the challenge of selecting meaningful augmentations on the fly, Luo et al. (2023) propose an adaptive data augmentation method, InfoTS, that uses information-aware criteria for selecting the optimal data augmentations to generate feasible positive samples for contrastive learning. The positive samples generated by InfoTS exhibit both high fidelity and diversity, which are essential for contrastive learning. Likewise, AutoTCL (Zheng et al., 2024) adaptively augments data based on a proposed input factorization technique to avoid ad-hoc decisions and reduce trial-and-error tuning by using a neural network to separate an instance into informative and task-irrelevant parts. The informative part is then transformed to preserve semantics, creating positive samples for learning. The method increases diversity in the augmented data and optimizes the augmentation process alongside contrastive learning. Li et al. (2024) introduce a novel approach to address the limitations of existing time-series models, which often suffer from high bias and low generalization due to predefined and rigid augmentation methods. The authors propose a unified and trainable augmentation operation that preserves patterns and minimizes bias by leveraging spectral information, which is designed to handle cross-domain datasets, making it highly scalable and generalizable.
For video data, Chen et al. (2022) use overlap augmentation techniques by sampling two subsequences with an overlap for each video. Overlapped timestamps are positives, and clips from other videos are negatives. Two timestamps neighboring each other also become positive pairs with Gaussian weight proportional to temporal distance. Zhang and Crandall (2022) present a novel technique that enhances the understanding of both spatial and temporal features. They achieve this by using temporal and spatial augmentations as a form of regularization, guiding the network to learn the desired semantics within a contrastive learning framework. Kim et al. (2023c) propose DynaAugment by dynamically changing the augmentation magnitude over time in order to learn the temporal variations of a video. By Fourier sampling the magnitudes, it regulates the augmentations diversely while maintaining temporal consistency. FreqAug (Kim et al., 2023b) pushes the model to focus more on dynamic features via dropping spatial or temporal low-frequency components. For various downstream tasks, FreqAug stochastically eliminates particular frequency components from the video to help the latent representation better capture meaningful features from the remaining data.
3.2.2. Sample Generation and Curation